A typical sugar cube has an edge length of 1 cm. If you had a cubical box that contained a mole of sugar cubes, what would it's edge length be?(1 mole =6.02×10^23

Grade:11

2 Answers

riya
7 years ago
Please answer it.I am not getting it.it is a  question which iam not able to solve.please hurry up.😩   kindly answer it fast.                                                                                                                                                              
                                              
Adarsh
7 years ago
The edge length and the corresponding number of cubes required for it are given below
1cm – 1
2cm – 8
3cm – 27
4cm –  64
and so on.
 
From this we get if ‘x’ is the number of cubes, then edge length = x1/3
So here, the reqiured edge length = (6.02×1023)1/3 = 8.44×107cm.
 
Hope you understand my solution.
Please approve only if my answer is correct.